OpenStudy (anonymous):

Express the ratio of the area of the larger circle to the area of the smaller circle in simplest radical form. Larger circle: radius of 2 + √3 and smaller circle's radius: 2-√3.

OpenStudy (anonymous):



OpenStudy (dumbcow):

is this a repost? what was unclear on explainations in prev post?

OpenStudy (anonymous):

Why is it 7 + 4√3?

OpenStudy (anonymous):

I tried pi (2)^2 + √3^2

OpenStudy (dumbcow):

oh you forgot the middle term (2+sqrt3)^2 = 4 + 2sqrt3 +2sqrt3 + 3

OpenStudy (dumbcow):

(a+b)^2 does Not equal (a^2 +b^2)

OpenStudy (anonymous):

middle term? what? D:

OpenStudy (dumbcow):

FOIL ? (2+sqrt3)(2+sqrt3) = 4+4sqrt3 +3

OpenStudy (anonymous):

ohhh.

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

Okay so I have now: \[\pi \times 7 + 4\sqrt{3} \div \pi \times 7-4\sqrt{3} \times \pi \times 7+4\sqrt{3} \div \pi \times 7+4\sqrt{3}\]

OpenStudy (dumbcow):

yes except the pi cancels out...pi/pi = 1 so you can get rid of it

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

Is that the same for 7/7?

OpenStudy (dumbcow):

no because its not a multiplicative constant...or its part of the addition so it can't cancel it helps to put parenthesis around them, (7+4sqrt3) is like 1 term so only if (7+4sqrt3) is on bottom can it cancel sorry if that was confusing

OpenStudy (anonymous):

Okay so: = \[(7 + 4\sqrt{3})(7 + 4\sqrt{3}) \div (7 - 4\sqrt{3})(7 + 4\sqrt{3})\] = \[49 + 28\sqrt{3} + 28\sqrt{3} + 16\sqrt{9} \div 49 + 28\sqrt{3} - 28\sqrt{3} - 16\sqrt{9}\] = \[121\sqrt{15} \div \]

OpenStudy (anonymous):

\[33\sqrt{15}\]

OpenStudy (anonymous):

The answer is supposed to be: 97 + 56√3

OpenStudy (dumbcow):

correct until last line...where did 121 come from?

OpenStudy (dumbcow):

also the numbers inside the radicals doNOt change when adding

OpenStudy (dumbcow):

oh i see what you did ... 2sqrt3 + 3sqrt2 does NOT equal 5sqrt5 keep the terms with radicals separate you can only combine sqrt2 with other sqrt2 's

OpenStudy (dumbcow):

Top: 16sqrt9 = 16*3 = 48 28sqrt3 + 28sqrt3 = 56sqrt3 49 + 48 = 97 --> 97 + 56sqrt3 Bottom: 49 - 48 = 1 28sqrt3 - 28sqrt3 = 0 --> 1

OpenStudy (anonymous):

ohhhh. Thank you!